# Orbits of particles with magnetic dipole moment around magnetized Schwarzschild black holes: Applications to S2 star orbit

### v1

@@ III.2 Periodic orbits for fixed topological characteristics and the impact of eccentricities
where $e$ and $\lambda$ denote eccentricity and the latus rectum, respectively. For bound orbits, the condition $10\leq e<$ 1 applies. Then we will have
@@
Figure 4: Orbits corresponding to various eccentricities: blue orbits represent $\beta=0.1$, while black orbits represent $\beta=0.0$. The black hole is assumed to be positioned at the origin.
@@ III.3 Estimating the magnetic coupling parameter $\beta$ using observed data from the S2 star
In this section, we will shift our focus from analyzing bound orbits to exploring unbound trajectories, particularly those involving deflection, which hold significant implications for understanding the dynamics of magnetized particles in black hole spacetimes.
@@ IV.2 OFK
From these diagrams, it is evident that particles exhibit two distinct turning point radii: the smallest orbital separation $r_{s}$ and the largest orbital separation $r_{l}$. The latter, $r_{l}$, corresponds to the OFK. Particles reaching this orbit either proceed to infinity following hyperbolic motion in the $\mathcal{E}^{2}_{u}>\mathcal{E}^{2}\geq 1-\beta$ case or engage in periodic orbits when $r_{l}=r_{a}$, as observed previously, in the $\mathcal{E}^{2}<1-\beta$ scenario. Notably, the equation of motion governing OFK is almost identical to that of bound orbits. However, it is crucial to consider that this is one of the roots of Eq. (18) is negative for unbound orbits. Consequently, assuming that a magnetized particle initiates its motion from $u=u_{2}$ and $\phi=0$, Eq. (21) can be expressed as
@@ IV.3 OSK
In Fig. 9, we illustrate the trajectories of magnetized particles approaching the black hole from a distance $r_{s}$, with fixed values of $\beta$ and the impact parameter $b={l}/{\mathcal{E}}$. From the figure, we infer that a larger impact parameter results in a more noticeable bending of the trajectories as particles approach the black hole, with the turning point positioned at a smaller distance. The magnetic interaction further enhances this bending, especially evident in the final segment before the particles fall into the black hole. Conversely, smaller impact parameters cause the trajectories to start their approach from a more distant location, indicating a turning point at a greater distance from the black hole.
@@ V Radial trajectories
Studying particles in free fall with zero angular momentum offers several advantages for standard general relativistic tests. One key area of investigation is the gravitational clock effect experienced by falling observers in gravitational fields, which forms the theoretical basis for phenomena such as the gravitational redshift-blueshift of light rays passing near a black hole. Another significant aspect is the phenomenon of frozen infalling objects, observed by distant observers as they approach the event horizon of a black hole. This stems from the disparities in time measurements between distant observers ($t$) and falling observers ($\tau$), and provides valuable insight into gravitational effects and their implications for our understanding of spacetime dynamics Zel’dovich et al. (1972); Ryder (2009); Zel’dovich and Novikov (2014). Hence, the radial motion of magnetized particles without angular momentum also presents an interesting subject to explore.
@@ VI Summary and conclusions
This study investigated particles’ bound and unbound orbits with a dipole moment as they approached a magnetized Schwarzschild black hole. Magnetization of the black hole was achieved by assuming proportionality between the four-potential and a space-like Killing vector. We constructed the Lagrangian dynamics for test particles, where the magnetic moment directly contributed to their conjugate momenta, with a magnetic interaction parameter $\beta$ playing a crucial role in determining the magnetic field’s effects on particle motion. We derived and analyzed the effective potential experienced by approaching particles, classifying possible bound and unbound orbits based on their specific energy and angular momentum. Bound orbits resembled planetary orbits oscillating between apoapsis and periapsis, with the possibility of complete circularity at the minimum of the effective potential, determining the ISCO. We noted that while the ISCO radius increased with $\beta$, it caused decreases in the specific energy and angular momentum of particles on the ISCO. Analytical solutions to the angular equations of motion for bound orbits were expressed in terms of elliptic integrals and Jacobi elliptic functions, with a precise classification of orbits based on topological characterization. Simulating various bound orbits with different eccentricities revealed topological variations and geometric extensions in the exterior spacetime of the black hole. Comparisons between observed positions of the S2 star around Sgr A* and theoretical predictions were made for two different values of $\beta$, calculating minimum and maximum values for this parameter, as well as specific energy and specific angular momentum of the S2 star. We found that the magnetic dipole moment associated with the S2 star was on the order of $10^{6}\,\mathrm{G\cdot cm^{3}}$. Unbound orbits, classified as unstable circular orbits, OFK, and OSK, either deflected from the black hole to infinity or fell inexorably onto the event horizon. The period of unstable circular orbits was calculated for comoving and distant observers, and the OFK was likened to the magnetic counterpart of gravitational Rutherford scattering. For such deflecting trajectories, the characteristic polynomial of the equation of motion had only two positive roots, leading to analytical solutions differing in elliptic integrals’ arguments. The scattering of test particles could be attractive or repulsive for the OFK, depending on eccentricity. The OSK trajectory curvature increased with the magnetic interaction parameter, resulting in a smaller turning point. Lastly, we explored purely radial orbits for magnetized particles, obtaining the radial profile of proper time for infalling observers and coordinate time for distant observers. The study of magnetized particles in black hole spacetimes offers new avenues for investigating relativistic stars at galaxy centers potentially associated with magnetic fields. Expanding models to include black holes with more degrees of freedom could advance research in relativistic astrophysics, a direction for future studies.
@@ Acknowledgements
$\Theta=\lambda^{2}M\Bigr{\{}\beta\Big{[}\beta\Big{(}6\beta(2\beta+3)-1\Big{)}+\Big{(}\beta\left(4\beta^{2}+6\beta-7\right)-5\Big{)}e^{2}-19\Big{]}-8\Bigr{\}}.$ (51d)

### v2

@@ III.2 Periodic orbits for fixed topological characteristics and the impact of eccentricities
where $e$ and $\lambda$ denote eccentricity and the latus rectum, respectively. For bound orbits, the condition $10\leq e<$ 1 applies. Then, we can have
@@
which possess the $\beta$-linearized forms
$\mathcal{E}^{2}=\frac{\Theta_{0}+\Phi_{0}}{\lambda^{2}\Psi_{0}},\quad l^{2}=\frac{\lambda M\Xi_{0}}{\Psi_{0}},$ (38)
@@
Figure 4: Orbits with fixed eccentricity for different taxonomy schemes. In the left panels, blue orbits correspond to $\beta=0.1$, while black orbits correspond to $\beta=0.0$. In the right panels, two different values of the gauge-invariant quantity, the impact parameter $b=\frac{l}{\mathcal{E}}$, are considered. The black hole is assumed to be located at the origin.
@@ III.3 Estimating the magnetic coupling parameter $\beta$ using observed data from the S2 star
In this section, we will shift our focus from analyzing bound orbits to exploring unbound trajectories, particularly those involving deflection. These have significant implications for understanding the dynamics of magnetized particles in black hole spacetimes.
@@ IV.2 OFK
These diagrams show that particles exhibit two distinct turning point radii: the smallest orbital separation $r_{s}$ and the largest orbital separation $r_{l}$. The latter, $r_{l}$, corresponds to the OFK. Particles reaching this orbit either proceed to infinity following hyperbolic motion in the $\mathcal{E}^{2}_{u}>\mathcal{E}^{2}\geq 1-\beta$ case or engage in periodic orbits when $r_{l}=r_{a}$, as observed previously, in the $\mathcal{E}^{2}<1-\beta$ scenario. Notably, the equation of motion governing OFK is almost identical to that of bound orbits. However, it is crucial to consider that this is one of the roots of Eq. (20) is negative for unbound orbits. Consequently, assuming that a magnetized particle initiates its motion from $u=u_{2}$ and $\phi=0$, Eq. (25) can be expressed as
@@ IV.3 OSK
In Fig. 9, we illustrate the trajectories of magnetized particles approaching the black hole from a distance $r_{s}$, with fixed values of $\beta$ and the impact parameter $b={l}/{\mathcal{E}}$. From the figure, we infer that a larger impact parameter results in a more noticeable bending of the trajectories as particles approach the black hole, with the turning point positioned at a smaller distance. The magnetic interaction further enhances this bending, especially in the final segment before the particles fall into the black hole. Conversely, smaller impact parameters cause the trajectories to start their approach from a more distant location, indicating a turning point at a greater distance from the black hole.
@@ V Radial trajectories
Studying particles in free fall with zero angular momentum offers several advantages for standard general relativistic tests. One key area of investigation is the gravitational clock effect experienced by falling observers in gravitational fields. This forms the theoretical basis for phenomena such as the gravitational redshift-blueshift of light rays passing near a black hole. Another significant aspect is the phenomenon of frozen infalling objects, observed by distant observers as they approach the event horizon of a black hole. This stems from the disparities in time measurements between distant observers ($t$) and falling observers ($\tau$), and provides valuable insight into gravitational effects and their implications for our understanding of spacetime dynamics Zel’dovich et al. (1972); Ryder (2009); Zel’dovich and Novikov (2014). Hence, the radial motion of magnetized particles without angular momentum also presents an interesting subject to explore.
@@ VI Summary and conclusions
This study investigated particles’ bound and unbound orbits with a dipole moment as they approached a magnetized Schwarzschild black hole. Magnetization of the black hole was achieved by assuming proportionality between the four-potential and a space-like Killing vector. We constructed the Lagrangian dynamics for test particles, where the magnetic moment directly contributed to their conjugate momenta, with a magnetic interaction parameter $\beta$ playing a crucial role in determining the magnetic field’s effects on particle motion. We derived and analyzed the effective potential experienced by approaching particles, classifying possible bound and unbound orbits based on their specific energy and angular momentum. Bound orbits resembled planetary orbits oscillating between apoapsis and periapsis, with the possibility of complete circularity at the minimum of the effective potential, determining the ISCO. We noted that while the ISCO radius increased with $\beta$, it caused decreases in the specific energy and angular momentum of particles on the ISCO. Analytical solutions to the angular equations of motion for bound orbits were expressed in terms of elliptic integrals and Jacobi elliptic functions, with a precise classification of orbits based on topological characterization. Simulating various bound orbits with different eccentricities revealed topological variations and geometric extensions in the exterior spacetime of the black hole. Comparisons between observed positions of the S2 star around Sgr A* and theoretical predictions were made for two different values of $\beta$, calculating minimum and maximum values for this parameter and specific energy and specific angular momentum of the S2 star. We found that the magnetic dipole moment associated with the S2 star was on the order of $10^{6}\,\mathrm{G\cdot cm^{3}}$. Unbound orbits, classified as unstable circular orbits, OFK, and OSK, either deflected from the black hole to infinity or fell inexorably onto the event horizon. The period of unstable circular orbits was calculated for comoving and distant observers, and the OFK was likened to the magnetic counterpart of gravitational Rutherford scattering. For such deflecting trajectories, the characteristic polynomial of the equation of motion had only two positive roots, leading to analytical solutions differing in elliptic integrals’ arguments. The scattering of test particles could be attractive or repulsive for the OFK, depending on eccentricity. The OSK trajectory curvature increased with the magnetic interaction parameter, resulting in a more minor turning point. Lastly, we explored purely radial orbits for magnetized particles, obtaining the radial profile of proper time for infalling and coordinating time for distant observers. Studying magnetized particles in black hole spacetimes offers new avenues for investigating relativistic stars at galaxy centers potentially associated with magnetic fields. Expanding models to include black holes with more degrees of freedom could advance research in relativistic astrophysics, a direction for future studies.
@@ Acknowledgements
$\Theta=\lambda^{2}M\Bigr{\{}\beta\Big{[}\beta\Big{(}6\beta(2\beta+3)-1\Big{)}+\Big{(}\beta\left(4\beta^{2}+6\beta-7\right)-5\Big{)}e^{2}-19\Big{]}-8\Bigr{\}},$ (57d)
$\Psi_{0}=2\left[\lambda-\left(e^{2}+3\right)M\right]^{2},$ (57e)
$\Xi_{0}=(2-3\beta)\lambda^{2}+4M^{2}\beta\left(1+3e^{2}\right)+\lambda M\left[19\beta+(5\beta-2)e^{2}-6\right],$ (57f)
$\Theta_{0}=-2(\beta-1)\lambda^{4}+4\beta\left(e^{2}-5\right)\left(e^{2}-1\right)^{2}M^{4}+8\left(e^{2}-1\right)\lambda M^{3}\left(-6\beta+e^{2}+3\right)$
$\qquad\qquad\qquad+\lambda^{3}M\left[15\beta+(\beta-2)e^{2}-14\right],$ (57g)
$\Phi_{0}=\lambda^{2}M^{2}\left(\beta\left(3e^{4}+2e^{2}-37\right)+32\right).$ (57h)